16× e htps/ html Joey Logano won the 57th Daytona 500 in February. If he pa loan (as id back a $270 loan with $630 interest at 70%, what wasthe time of the sume 360 days)? (Do not round intermediate calculations. Round up your answer to the nearest day.) of the
Answer #1

We use the below formula to calculate the time of the loan

Interest = Principal * time * interest rate

6.30 = 270 * No. of days/360 * 0.07

6.30 = 18.9* No. of days/360

No. of days = 6.3*360/18.9

Number of days = 120

Time of the loan = 120 days